OpenStudy (anonymous):

If (a + 3) / 5 is an integer, what is remainder when ‘a’ is divided by 5? A. 1 B. 2 C. 3 D. 4 E. 5

OpenStudy (mathmath333):

sry 2

OpenStudy (anonymous):

Sorry. But can u explain how? ^^'

OpenStudy (mathmath333):

yes

OpenStudy (mathmath333):

for an integer the remainder of (a+3)/5 must be zero

OpenStudy (mathmath333):

possible values of a=2,7,12,17..........

OpenStudy (mathmath333):

dividend =quotent *divisor +remainder

OpenStudy (mathmath333):

(a+3)=q*5+r

OpenStudy (mathmath333):

but r=0

OpenStudy (mathmath333):

a+3=q*5

OpenStudy (mathmath333):

a=q*5-3

OpenStudy (mathmath333):

now substitute q =1,2,3,4,5.... u will get a=2,7,12,17....

OpenStudy (mathmath333):

hence the answer is 2

OpenStudy (anonymous):

Sorry Brother. I still dont quite get it. To be honest.. I dont even know what they are asking for in the question.

ganeshie8 (ganeshie8):

a+3 = 5k a = -3 + 5k = (2-5) + 5k = 2 +5(k-1) = 2 + 5m

ganeshie8 (ganeshie8):

so "a" leaves a remainder "2" when divided by 5

OpenStudy (mathmath333):

u know what integer means?

OpenStudy (anonymous):

To be Honest. Nope @mathmath333 . And what is k? @geneshie8

OpenStudy (mathmath333):

well integer means a number which has no denominator ,like ...-3,-2,-1,0,1,2,3.......

OpenStudy (mathmath333):

where as 1/2 ,2/3 ,1/3 are not integers

OpenStudy (mathmath333):

1/2=0.5 it leaves decimal

OpenStudy (anonymous):

I see.. I get it.

OpenStudy (mathmath333):

hence for the result of (a+3)/5 to be integer it should not have denominator that is it should not have any decimal

OpenStudy (anonymous):

I see..

OpenStudy (mathmath333):

hope u got it

OpenStudy (anonymous):

More or less..

OpenStudy (anonymous):

"what is remainder when a is divided by 5 ?" What does this line means?

OpenStudy (anonymous):

Oh well.. Thanks anyway brother.

OpenStudy (mathmath333):

for ex if 4 is divided by 2 the remainder is 0

OpenStudy (anonymous):

I guess it is.